OpenStudy (anonymous):

Find the product of: (x-3)(x^2+3x+9)

OpenStudy (anonymous):

\[x^3-27\]

OpenStudy (anonymous):

I think that might be right

OpenStudy (anonymous):

wow thank you

OpenStudy (anonymous):

\[(x^{3}+3x^{2}+9x)+ (-3x^{2}-9x-27) = x^{3}-27\]

OpenStudy (anonymous):

my pleasure

OpenStudy (anonymous):

well that is a simple formula \[x^3 - y^3\] = \[x-y\] .\[x^2 + xy + y^2\] this is what i did here x = x and y = 3

OpenStudy (anonymous):

@rayray go back to this one and see the changes http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4e75786d0b8b7d4f6cfc8ee6